# Critical behaviour of the fully packed loop-$O(n)$ model on planar triangulations

### v2

@@ 2.1.1 Definition of the model
The weight depends only on the loop ensemble $L(\mathfrak{m},\omega)$ and therefore is independent of the choice of root edge; in other words, the root of $M$ is sampled uniformly once the map is fixed. Conditionally on $M$, the edge configuration $\Omega$ is exactly the self-dual FK(q)(q) (or random-cluster) measure on $M$; see, for instance, [duminil2013parafermionic].
@@ 2.1.2 The Mullin–Bernardi–Sheffield bijection
We summarise here the bijective constructions of Mullin [mullin67], Bernardi [bernardi08] and Sheffield [sheffield2016quantum]; see also [BP, Chapter 4] for another presentation of these ideas. These constructions relate planar maps equipped with a distinguished subset of edges to words written in the alphabet $\Theta$ introduced below. The latter may be interpreted as trajectories of a two-type inventory system evolving in discrete time.
@@
One checks [sheffield2016quantum, Section 4.1] that companion triangles always form matching burger–order pairs under the reduction rules; in particular, this implies that every burger created from a triangle lying inside one of the original loops is consumed before the exploration leaves this loop. Therefore, we have $\overline{w}=\emptyset$. Importantly, we note that the number of $\mathsf{F}$ symbols is exactly one fewer than the number of loops in $L(\mathfrak{m},\omega)$.
@@ 2.2.1 Definition of the model
Recall from the preceding section that a finite FK(q)(q)-decorated planar map $(M,\Omega)=(M_{k},\Omega_{k})$ with $k$ edges and cluster-weight parameter $q$ may be encoded by a word $W$ sampled under law (2.4). The correspondence between the map and the word is provided by the Mullin–Bernardi–Sheffield bijection presented in Section 2.1.2. As established in [chen2017basic] and [sheffield2016quantum], as $k\to\infty$ the sequence $(M_{k},\Omega_{k})$ converges in distribution, with respect to the local topology, to a limiting object $(M_{\infty},\Omega_{\infty})$ known as the infinite FK(q)(q) planar map. The metric governing this convergence can be taken to be
@@
First, we know from [sheffield2016quantum, Proposition 2.2] that every letter in the bi-infinite sequence (2.5) has a unique match: each burger symbol ($\mathsf{c}$ or $\mathsf{h}$) is eventually consumed by an order symbol ($\mathsf{C}$, $\mathsf{H}$, or $\mathsf{F}$), and every order corresponds to some earlier burger. The match of $X(i)$ is denoted $X(\varphi(i))$.
@@
One can iterate this over the full nested sequence of excursions containing $X(0)$ to get an exhaustion of the infinite triangulation by growing neighbourhoods of the root. For each graph-distance radius $R$, the associated ball is described by the unique excursion whose envelope first contains all vertices within that distance. Since these maps are consistent with one another, the infinite object $(M_{\infty},\Omega_{\infty})$ is defined as a “projective limit”. An alternative concrete and elegant description of $(M_{\infty},\Omega_{\infty})$ is given in [chen2017basic].
@@ 2.2.2 The reduced burger walk
We recall the construction of the reduced burger walk of [berestycki2017critical], which is the key exploration that we will use to derive our loop and cluster exponents in 1.3. First, we need to introduce some terminology on hamburger-cheeseburger words. Fix a bi-infinite i.i.d. sequence $W=(X_{n})_{n\in\mathbb{Z}}$ as in (2.5). Recall that a word $e$ is called an $\mathsf{F}$-excursion (of $W$) if it is of the form $\mathsf{h}\cdots\mathsf{F}$ (type $\mathsf{h}$) or $\mathsf{c}\cdots\mathsf{F}$ (type $\mathsf{c}$), where the final $\mathsf{F}$ is matched to the first letter. We say that an $\mathsf{F}$-excursion which is subword of $(X_{n})_{n\leq 0}$ is maximal if it is not contained in any other $\mathsf{F}$-excursion inside $X(-\infty,0)$. Then we can write $X(-\infty,0)$ in a unique way as
@@
Note that the increments of the reduced walk are always one-dimensional. Specifically, $\tilde{h}$ can only move at times corresponding to words in $\mathscr{A}_{\mathsf{h}}$, while $\tilde{c}$ can only move at times corresponding to words in $\mathscr{A}_{\mathsf{c}}$. For this reason, it will often be more convenient to work with the walk $(h_{n},n\geq 0)$ (resp. $(c_{n},n\geq 0)$) skipping past times when $h$ (resp. $c$) stays put. Let $\tau^{\mathsf{h}}$ (resp. $\tau^{\mathsf{c}}$) the hitting times of 1-1 by $h$ (resp. $c$). By symmetry between ham and cheeseburgers, the two random walks $h$ and $c$ have the same step distribution, hence $\tau^{\mathsf{h}}$ and $\tau^{\mathsf{c}}$ have the same distribution. Moreover, a key observation of [berestycki2017critical] is that $h$ and $c$ are actually independent, by independence of the symbols in the hamburger-cheeseburger sequence, and the fact that the $\mathsf{F}$-excursions are processed all at once.
@@
In particular, we claim that the random walks $h$ and $c$ are centred. Indeed, [sheffield2016quantum, Section 3.1] implies that the variable $\Xi$ has expectation $1$: in Sheffield’s notation, $\Xi$ is nothing but the law of |X(J,1)|1|X(-J,-1)|-1 (conditional on $X(0)=\mathsf{F}$, which is independent), so that $\mathbb{E}[\Xi]=\chi-1=1$. Therefore
@@ 2.3.1 Definition and connection with FK-decorated maps
It is convenient to interpret this model as the symmetric specialisation of the twofold loop model studied in [BBG12]. In that formulation, the loops divide the triangulation into regions coloured red or blue, the colouring of each triangle depending solely on the parity of the number of loops encircling that triangle. Because no loop crosses the root face, the whole boundary loop (or rather the triangles traversed by it) is monochromatic; thus the entire colouring is fixed once the boundary colour is prescribed.
@@
where $\mathfrak{f}_{1}$ and $\mathfrak{f}_{2}$ correspond to the triangles of colour 1 (blue) and 2 (red) of $\mathfrak{t}$ respectively, and $x_{1},x_{2}>0$.
@@
We now recall the link with FK-decorated maps. Implicit in the work of [BBG12] is the fact that the Tutte map $T(\mathfrak{m},\omega)$ of a self-dual FK(q)(q)-weighted map $(\mathfrak{m},\omega)$ with $q=q(p)$ (see Section 2.1.1) has the law of a fully packed loop-$O(n)$ triangulation provided that the paramaters are suitably chosen, which turns out to require the following relations:
@@ 2.3.2 The gasket decomposition
The gasket decomposition, introduced in [BBG12a] and further developed in [BBG12], provides a remarkably effective framework for analysing a broad class of loop models. In what follows we adapt the presentation of [BBG12] to the particular fully packed setting determined by (2.11) and (2.12).
Consider a configuration $(\mathfrak{t},\boldsymbol{\ell})\in\mathbb{T}_{\ell}$. To construct its gasket, we first look at the edges of $\mathfrak{t}$ that are reachable from the boundary without crossing any loop. Because the loops are fully packed, these accessible edges divide the map into exactly $\boldsymbol{\ell}+1$ faces: the external face, with boundary length $\ell$, and an additional face of degree $k$ for each loop of (outer) perimeter $k$. The resulting object is by definition the gasket. Now examine one of the internal faces of degree $k$ created by the gasket.
@@
If we reinsert the triangles traversed by the corresponding loop, we obtain a ring of triangles whose outer boundary length is $k$ and whose inner boundary has some length $k^{\prime}$. The ring consists of $k+k^{\prime}$ triangles in total. Moreover, the loop-decorated triangulation originally contained inside this loop – call it $(\mathfrak{t}^{\prime},\boldsymbol{\ell}^{\prime})\in\mathbb{T}_{k^{\prime}}$ – is precisely what was taken out from $(\mathfrak{t},\boldsymbol{\ell})$ to form the corresponding face of the gasket. In this way, $(\mathfrak{t},\boldsymbol{\ell})$ decomposes into: (i) the gasket; (ii) one ring of triangles for each internal gasket face; and (iii) a fully packed loop-decorated triangulation attached along the inner boundary of each such ring. The contribution of such a component to the weight $Z(\mathfrak{t}^{\prime},\boldsymbol{\ell}^{\prime},x,n)$ in (2.11) is $nx^{k+k^{\prime}}Z(\mathfrak{t},\boldsymbol{\ell},x,n)$, where $k$ is the outer and $k^{\prime}$ the inner boundary length of the ring. This motivates the definition
$g_{k}:=n\sum_{k^{\prime}=0}^{\infty}A_{k\rightarrow k^{\prime}}x^{k+k^{\prime}}F_{k^{\prime}},$ (2.16)
@@
where $A_{k\rightarrow k^{\prime}}$ counts the possible rings of triangles with outer boundary $k$ and inner boundary $k^{\prime}$. See Figure 3. Thus $g_{k}$ corresponds to the weight of any face of degree $k$ in the gasket of $(\mathfrak{t},\boldsymbol{\ell})$. In other words, this gasket has the law of a Boltzmann map with fixed boundary length $\ell\geq 1$, and where we assign to each face of degree $m$ the weight $g_{m}$. We denote by $\mathcal{F}_{\ell}((g_{m}){m\geq 1})$ the resulting partition function, and note that this coincides with $F_{\ell}$. Thus (2.16) may be rewritten as the so-called fixed point equation
$g_{k}=n\sum_{k^{\prime}=0}^{\infty}A_{k\rightarrow k^{\prime}}x^{k+k^{\prime}}\mathcal{F}_{k^{\prime}}((g_{m})_{m\geq 1}).$ (2.17)
Introduce the resolvent
$W(z)=\sum_{\ell=0}^{\infty}\frac{\mathcal{F}_{\ell}((g_{m})_{m\geq 1})}{z^{\ell+1}}=\sum_{\ell=0}^{\infty}\frac{F_{\ell}}{z^{\ell+1}},$ (2.18)
@@
which plays a crucial role in what follows. It is known that such a solution satisfies the one-cut lemma [BBG12b]. More precisely, it is known that $W$ is analytic on $\mathbb{C}\setminus[\gamma_{-},\gamma_{+}]$ for some real interval $\gamma_{-},\gamma_{+}]$ containing zero, $\gamma_{-}\leq 0\leq\gamma_{+}$, with $\gamma_{+}\geq|\gamma_{-}|$, and that it satisfies $W(z)\sim 1/z$ as $|z|\to\infty$. Moreover, the spectral density $\rho$ defined by
$\rho(y):=-\frac{W(y+i0)-W(y-i0)}{2\pi i},\quad y\in[\gamma_{-},\gamma_{+}],$ (2.19)
@@
is positive in $(\gamma_{-},\gamma_{+})$ and continuous on $\gamma_{-},\gamma_{+}]$, which vanishes at both endpoints. These properties have been established rigorously using combinatorial bijections with Motzkin paths, see [BBG12b, Section 6]. By Cauchy integration, the spectral density determines the resolvent via the following formula:
$W(z)=\int^{\gamma_{+}}_{\gamma_{-}}\frac{\rho(y)}{z-y}\mathrm{d}y,\quad z\in\mathbb{C}\setminus[\gamma_{-},\gamma_{+}].$
With this notation, equations (3.22)–(3.23) of [BBG12] specialise to the relation
@@
which is called the resolvent equation. We note that we also have $\gamma_{+}\leq 1/(2x)$, in order for the equation to actually make sense: indeed, note that we need the argument $1/x-z\notin(\gamma_{-},\gamma_{+})$ for the left hand side to be meaningful. This imposes some requirements: the image of the interval $(\gamma_{-},\gamma_{+})$ by $z\mapsto 1/x-z$ must lie entirely outside of $(\gamma_{-},\gamma_{+})$ and so must lie entirely either one the left or on the right of this interval. For $z=1/x-\gamma_{-}>0\geq\gamma_{-}$ so that in fact it must lie entirely on the right of the interval. In other words $1/x-z\geq\gamma_{+}$ for all $z\in(\gamma_{-},\gamma_{+})$. Choosing $z=\gamma_{+}-\epsilon$ and letting $\epsilon\to 0$ we deduce that
@@
This fact is stated just above (3.19) in [BBG12]. When $x=x_{c}$ we will see another argument based on the hamburger–cheeseburger bijection below (see Proposition 3.6, in particular (3.9)). One particular goal of the present paper is to solve this equation explicitly in the case when $n\in(0,2)$ and $x=x_{c}$ as in (2.14).
The works [BBG12a, BBG12b, BBG12] solve the resolvent equation (2.20) when the cut $\gamma_{-},\gamma_{+}]$ is fixed, giving an explicit expression in the elliptic parametrisation (in a far more general framework than fully packed triangulations). However, equation (2.20) should be thought of as an equation on the triplet $(W,\gamma_{-},\gamma_{+})$, where the cut is part of the unknown. The aforementioned works relied on numerical evidence to support this uniqueness ansatz (i.e., there is really a unique choice of $\gamma_{-},\gamma_{+}$, and therefore ultimately of $W$). In the present paper, we will first derive some information about the cut (Section 3), before we solve the equation. Because of our more restrictive framework of fully packed triangulations, our solution in Section 4 will actually be closer to the physics paper of Gaudin and Kostov [gaudin1989n]. We refer to [BBD16, korzhenkova2022exploration] for a more detailed discussion.
We conclude by mentioning that this uniqueness ansatz was rigorously established by Budd and Chen [budd2019peeling] in the case of rigid quadrangulations. This approach uses the symmetry relation $\gamma_{-}=-\gamma_{+}$ that holds in the case of bipartite maps and which does not appear to admit a straightforward extension to our setting.

### v3

@@ 2.1.1 Definition of the model
The weight depends only on the loop ensemble $L(\mathfrak{m},\omega)$ and therefore is independent of the choice of root edge; in other words, the root of $M$ is sampled uniformly once the map is fixed. Conditionally on $M$, the edge configuration $\Omega$ is exactly the self-dual FK(q)(q) (or random-cluster) measure on $M$; see, for instance, [DUM13].
@@ 2.1.2 The Mullin–Bernardi–Sheffield bijection
We summarise here the bijective constructions of Mullin [MUL67], Bernardi [BER08] and Sheffield [SHE16]; see also [BP24, Chapter 4] for another presentation of these ideas. These constructions relate planar maps equipped with a distinguished subset of edges to words written in the alphabet $\Theta$ introduced below. The latter may be interpreted as trajectories of a two-type inventory system evolving in discrete time.
@@
One checks [SHE16, Section 4.1] that companion triangles always form matching burger–order pairs under the reduction rules; in particular, this implies that every burger created from a triangle lying inside one of the original loops is consumed before the exploration leaves this loop. Therefore, we have $\overline{w}=\emptyset$. Importantly, we note that the number of $\mathsf{F}$ symbols is exactly one fewer than the number of loops in $L(\mathfrak{m},\omega)$.
@@ 2.2.1 Definition of the model
Recall from the preceding section that a finite self-dual FK(q)(q)-decorated planar map $(M,\Omega)=(M_{k},\Omega_{k})$ with $k$ edges and cluster-weight parameter $q$ may be encoded by a word $W$ sampled under law (2.4). The correspondence between the map and the word is provided by the Mullin–Bernardi–Sheffield bijection presented in Section 2.1.2. As established in [CHE17] and [SHE16], as $k\to\infty$ the sequence $(M_{k},\Omega_{k})$ converges in distribution, with respect to the local topology, to a limiting object $(M_{\infty},\Omega_{\infty})$ known as the infinite (self-dual) FK(q)(q) planar map. The metric governing this convergence can be taken to be
@@
First, we know from [SHE16, Proposition 2.2] that every letter in the bi-infinite sequence (2.5) has a unique match: each burger symbol ($\mathsf{c}$ or $\mathsf{h}$) is eventually consumed by an order symbol ($\mathsf{C}$, $\mathsf{H}$, or $\mathsf{F}$), and every order corresponds to some earlier burger. The match of $X(i)$ is denoted $X(\varphi(i))$.
@@
One can iterate this over the full nested sequence of excursions containing $X(0)$ to get an exhaustion of the infinite triangulation by growing neighbourhoods of the root. For each graph-distance radius $R$, the associated ball is described by the unique excursion whose envelope first contains all vertices within that distance. Since these maps are consistent with one another, the infinite object $(M_{\infty},\Omega_{\infty})$ is defined as a “projective limit”. An alternative concrete and elegant description of $(M_{\infty},\Omega_{\infty})$ is given in [CHE17].
@@ 2.2.2 The reduced burger walk
We recall the construction of the reduced burger walk of [BLR17], which is the key exploration that we will use to derive our loop and cluster exponents in 1.3. First, we need to introduce some terminology on hamburger-cheeseburger words. Fix a bi-infinite i.i.d. sequence $W=(X_{n})_{n\in\mathbb{Z}}$ as in (2.5). Recall that a word $e$ is called an $\mathsf{F}$-excursion (of $W$) if it is of the form $\mathsf{h}\cdots\mathsf{F}$ (type $\mathsf{h}$) or $\mathsf{c}\cdots\mathsf{F}$ (type $\mathsf{c}$), where the final $\mathsf{F}$ is matched to the first letter. We say that an $\mathsf{F}$-excursion which is subword of $(X_{n})_{n\leq 0}$ is maximal if it is not contained in any other $\mathsf{F}$-excursion inside $X(-\infty,0)$. Then we can write $X(-\infty,0)$ in a unique way as
@@
Note that the increments of the reduced walk are always one-dimensional. Specifically, $\tilde{h}$ can only move at times corresponding to words in $\mathscr{A}_{\mathsf{h}}$, while $\tilde{c}$ can only move at times corresponding to words in $\mathscr{A}_{\mathsf{c}}$. For this reason, it will often be more convenient to work with the walk $(h_{n},n\geq 0)$ (resp. $(c_{n},n\geq 0)$) skipping past times when $h$ (resp. $c$) stays put. Let $\tau^{\mathsf{h}}$ (resp. $\tau^{\mathsf{c}}$) the hitting times of 1-1 by $h$ (resp. $c$). By symmetry between ham and cheeseburgers, the two random walks $h$ and $c$ have the same step distribution, hence $\tau^{\mathsf{h}}$ and $\tau^{\mathsf{c}}$ have the same distribution. Moreover, a key observation of [BLR17] is that $h$ and $c$ are actually independent, by independence of the symbols in the hamburger-cheeseburger sequence, and the fact that the $\mathsf{F}$-excursions are processed all at once.
@@
In particular, we claim that the random walks $h$ and $c$ are centred. Indeed, [SHE16, Section 3.1] implies that the variable $\Xi$ has expectation $1$: in Sheffield’s notation, $\Xi$ is nothing but the law of |X(J,1)|1|X(-J,-1)|-1 (conditional on $X(0)=\mathsf{F}$, which is independent), so that $\mathbb{E}[\Xi]=\chi-1=1$. Therefore
@@ 2.3.1 Definition and connection with FK-decorated maps
It is convenient to interpret this model as the symmetric specialisation of the twofold loop model studied in [BBG12b]. In that formulation, the loops divide the triangulation into regions coloured red or blue, the colouring of each triangle depending solely on the parity of the number of loops encircling that triangle. Because no loop crosses the root face, the whole boundary loop (or rather the triangles traversed by it) is monochromatic; thus the entire colouring is fixed once the boundary colour is prescribed.
@@
where $\mathfrak{f}_{1}$ and $\mathfrak{f}_{2}$ correspond to the (internal) triangle faces of colour 1 (blue) and 2 (red) of $\mathfrak{t}$ respectively, and $x_{1},x_{2}>0$. We will refer to the above weight as $Z(\mathfrak{t},\boldsymbol{\ell},x_{i},x_{j},n):=x_{1}^{\#\mathfrak{f}_{1}}x_{2}^{\#\mathfrak{f}_{2}}n^{\#\boldsymbol{\ell}}$.
@@
We now recall the link with FK-decorated maps. Implicit in the work of [BBG12b] is the fact that the Tutte map $T(\mathfrak{m},\omega)$ of a self-dual FK(q)(q)-weighted map $(\mathfrak{m},\omega)$ with $q=q(p)$ (see Section 2.1.1) has the law of a fully packed loop-$O(n)$ triangulation provided that the parameters are suitably chosen, which turns out to require the following relations:
@@
In the general bicoloured model, this transformation forms a one-to-one correspondence between the FK($q$) model of (1.1) and the loop-$O(n)$ model (2.13) with loop weight $n=\sqrt{q}$ and edge weights
$x_{1}:=s\Big(\frac{p_{0}}{\sqrt{q}(1-p_{0})}\Big)^{1/2}\quad\text{and}\quad x_{2}:=s.$
See [BBG12b, Section 2.1]. In particular, we see that pairs $(s,p_{0})$ in the region $\mathcal{R}_{q}$ defined in (1.9) simply correspond to taking $x_{1},x_{2}\leq x_{c}$. In particular, our 1.1 will entail that the corresponding partition function at $(s,p_{0})$ is finite.
@@ 2.3.2 The gasket decomposition
The gasket decomposition, introduced in [BBG12a] and further developed in [BBG12b], provides a remarkably effective framework for analysing a broad class of loop models. In what follows we adapt the presentation of [BBG12b] to the particular fully packed setting determined by (2.11) and (2.12).
Consider a configuration $(\mathfrak{t},\boldsymbol{\ell})\in\mathbb{T}_{\ell}$. We may see $(\mathfrak{t},\boldsymbol{\ell})$ as a bicoloured triangulation by colouring the boundary $(\mathfrak{t},\boldsymbol{\ell})$ with colour $i=1,2$ and alternating colours when crossing the loops. We denote by $j$ the other colour. To construct its gasket, we first look at the edges of $\mathfrak{t}$ that are reachable from the boundary without crossing any loop. Because the loops are fully packed, these accessible edges divide the map into exactly $\boldsymbol{\ell}+1$ faces: the external face, with boundary length $\ell$, and an additional face of degree $k$ for each loop of (outer) perimeter $k$. Note that all these faces have colour $i$. The resulting object is by definition the gasket. Now examine one of the internal faces of degree $k$ created by the gasket.
@@
If we reinsert the triangles traversed by the corresponding loop, we obtain a ring made up of triangles with colour $i$ whose outer boundary length is $k$ and triangles with colour $j$ whose inner boundary has some length $k^{\prime}$. See Figure 3. The ring consists of $k+k^{\prime}$ triangles in total. Moreover, the loop-decorated triangulation originally contained inside this loop — call it $(\mathfrak{t}^{\prime},\boldsymbol{\ell}^{\prime})\in\mathbb{T}_{k^{\prime}}$ — is precisely what was taken out from $(\mathfrak{t},\boldsymbol{\ell})$ to form the corresponding face of the gasket. That said pair $(\mathfrak{t}^{\prime},\boldsymbol{\ell}^{\prime})$ naturally comes with the colour $j$ which is the colour of its boundary in the original map $(\mathfrak{t},\boldsymbol{\ell})$. In this way, $(\mathfrak{t},\boldsymbol{\ell})$ decomposes into: (i) the gasket; (ii) one ring of triangles for each internal gasket face; and (iii) a fully packed loop-decorated triangulation attached along the inner boundary of each such ring. The contribution of such a component to the total weight $Z(\mathfrak{t},\boldsymbol{\ell},x_{i},x_{j},n)$ in (2.11) is $nA^{(i\to j)}_{k,k^{\prime}}Z(\mathfrak{t}^{\prime},\boldsymbol{\ell}^{\prime},x_{i},x_{j},n)$, where $k$ is the outer and $k^{\prime}$ the inner boundary length of the ring, and $A^{(i\to j)}_{k,k^{\prime}}$ is the partition function of rings of triangles of outer boundary with colour $i$ and size $k$, and inner boundary of colour $j$ and size $k^{\prime}$. This motivates the definition
$g_{k}^{(i)}:=n\sum_{k^{\prime}=0}^{\infty}A^{(i\to j)}_{k,k^{\prime}}F_{k^{\prime}}^{(j)},$ (2.16)
@@
See Figure 3. Thus $g_{k}^{(i)}$ corresponds to the weight of any face of degree $k$ in the gasket of $(\mathfrak{t},\boldsymbol{\ell})$ with boundary colour $i$. In other words, this gasket has the law of a Boltzmann map with fixed boundary length $\ell\geq 1$, and where we assign to each face of degree $m$ the weight $g_{m}^{(i)}$. We denote by $\mathcal{F}_{\ell}((g_{m}^{(i)}){m\geq 1})$ the resulting partition function, and note that this coincides with $F_{\ell}^{(i)}$. Thus (2.16) may be rewritten as the so-called fixed point equation
$g_{k}^{(i)}=n\sum_{k^{\prime}=0}^{\infty}A^{(i\to j)}_{k\rightarrow k^{\prime}}\mathcal{F}_{k^{\prime}}((g_{m}^{(j)})_{m\geq 1}).$ (2.17)
Introduce the resolvents
$W^{(i)}(z)=\sum_{\ell=0}^{\infty}\frac{\mathcal{F}_{\ell}((g_{m}^{(i)})_{m\geq 1})}{z^{\ell+1}}=\sum_{\ell=0}^{\infty}\frac{F_{\ell}^{(i)}}{z^{\ell+1}},$ (2.18)
@@
which play a crucial role in what follows. It is known that such a solution satisfies the one-cut lemma [BBG12c]. More precisely, it is known that $W^{(i)}$ is analytic on $\mathbb{C}\setminus{[\gamma^{(i)}_{-},\gamma^{(i)}_{+}]}$ for some real interval $\gamma^{(i)}_{-},\gamma^{(i)}_{+}]$} containing zero, ${\gamma_{-}^{(i)}}\leq 0\leq{\gamma_{+}^{(i)}}$, with ${\gamma_{+}^{(i)}}\geq|{\gamma_{-}^{(i)}}|$, and that it satisfies ${W^{(i)}}(z)\sim 1/z$ as $|z|\to\infty$. Moreover, the spectral density $\rho^{(i)}$} defined by
${\rho^{(i)}}(y):=-\frac{{W^{(i)}}(y+i0)-{W^{(i)}}(y-i0)}{2\pi i},\quad y\in{[\gamma^{(i)}_{-},\gamma^{(i)}_{+}]},$ (2.19)
@@
is positive in $(\gamma^{(i)}_{-},\gamma^{(i)}_{+})$} and continuous on $\gamma^{(i)}_{-},\gamma^{(i)}_{+}]$}, which vanishes at both endpoints. These properties have been established rigorously using combinatorial bijections with Motzkin paths, see [BBG12c, Section 6]. By Cauchy integration, the spectral density determines the resolvent via the following formula:
$W^{(i)}(z)=\int^{\gamma^{(i)}_{+}}_{\gamma^{(i)}_{-}}\frac{\rho^{(i)}(y)}{z-y}\mathrm{d}y,\quad z\in\mathbb{C}\setminus[\gamma^{(i)}_{-},\gamma^{(i)}_{+}].$
With this notation, equations (3.22)–(3.23) of [BBG12b] specialise to the two resolvent equations
$W^{(1)}(z+i0)+W^{(1)}(z-i0)+nW^{(2)}\left(s_{1}(z)\right)=z,\qquad z\in(\gamma^{(1)}_{-},\gamma^{(1)}_{+}),$ (2.20)
and
$W^{(2)}(z+i0)+W^{(2)}(z-i0)+nW^{(1)}\left(s_{2}(z)\right)=z,\qquad z\in(\gamma^{(2)}_{-},\gamma^{(2)}_{+}),$ (2.21)
where $s_{i}(z):=\frac{1}{x_{j}}(1-x_{i}z)$ for $i=1,2$. Note that for the above two equations (2.20)–(2.21) to make sense, we need $s_{1}(z)\notin(\gamma^{(2)}_{-},\gamma^{(2)}_{+})$ whenever $z\in(\gamma^{(1)}_{-},\gamma^{(1)}_{+})$. See the discussion right above equation (3.19) in [BBG12b]. Moreover, $s_{1}$ is continuous, so $s_{1}(z)$ has to stay on the same side of $(\gamma^{(2)}_{-},\gamma^{(2)}_{+})$ for all $z\in(\gamma^{(1)}_{-},\gamma^{(1)}_{+})$. Since $0\in(\gamma^{(1)}_{-},\gamma^{(1)}_{+})$ and $s_{1}(0)$ is clearly positive, it must hold that $s_{1}(z)>\gamma^{(2)}_{+}$ for all $z\in(\gamma^{(1)}_{-},\gamma^{(1)}_{+})$. This implies $\frac{1-x_{1}\gamma_{+}^{(1)}}{x_{2}}\geq\gamma^{(2)}_{+}$, which yields the inequality
$x_{1}\gamma_{+}^{(1)}+x_{2}\gamma_{+}^{(2)}\leq 1.$ (2.22)
In the colour-symmetric case (i.e. $x_{1}=x_{2}=x$), we will drop all superscripts $i$ and $j$ from the notation. Note that the two equations (2.20)–(2.21) then boil down to the single equation
@@
and that (2.22) yields
@@
When $x=x_{c}$ we will see another argument based on the hamburger–cheeseburger bijection below (see Proposition 3.6, in particular (3.9)). One particular goal of the present paper is to solve this equation explicitly in the case when $n\in(0,2)$ and $x=x_{c}$ as in (2.14).
The works [BBG12a, BBG12c, BBG12b] solve the resolvent equation (2.23) when the cut $\gamma_{-},\gamma_{+}]$ is fixed, giving an explicit expression in the elliptic parametrisation (in a far more general framework than fully packed triangulations). However, equation (2.23) should be thought of as an equation on the triplet $(W,\gamma_{-},\gamma_{+})$, where the cut is part of the unknown. The aforementioned works relied on numerical evidence to support this uniqueness ansatz (i.e., there is really a unique choice of $\gamma_{-},\gamma_{+}$, and therefore ultimately of $W$). In the present paper, we will first derive some information about the cut (Section 3), before we solve the equation. Because of our more restrictive framework of fully packed triangulations, our solution in Section 4 will actually be closer to the physics paper of Gaudin and Kostov [GK89]. We refer to [BBD23, KOR22] for a more detailed discussion.
We conclude by mentioning that this uniqueness ansatz was rigorously established by Budd and Chen [BC19] in the case of rigid quadrangulations. This approach uses the symmetry relation $\gamma_{-}=-\gamma_{+}$ that holds in the case of bipartite maps and which does not appear to admit a straightforward extension to our setting.